# A measurement-based protocol for the generation of delocalised quantum states of a mechanical system

### v1

@@ Abstract
Non-Gaussian mechanical states are a key resource for quantum-enhanced sensing and tests of macroscopic quantum physics. We propose a measurement-based protocol to herald delocalized, nonclassical states of a mechanical oscillator in cavity optomechanics by conditioning on Geiger photodetection of the optical output. We analyse under which conditions Stokes-induced optomechanical entanglement give rise to mechanical Wigner Function negativity upon detection. We develop and compare a blue-detuned pulsed scheme and a continuous-wave steady-state scheme employing temporal-mode filtering, and we quantify heralding rates and robustness to finite temperature under realistic detection efficiencies.
@@ I Introduction
Quantum optomechanics offers a possible route to macroscopic quantum control: massive oscillators comprising billions of atoms can be cooled, entangled, and measured at or near the quantum limit. A challenge is the generation of delocalized coherent mechanical states. Previous strategies implemented in levitated and clamped-cavity platforms typically introduce a nonlinearity or conditional operation to evolve semiclassical inputs into nonclassical outputs; examples include transient potentials, photon addition or subtraction, and conditional measurements [45, 34, 30, 41, 25, 52, 13] or the intrinsic optomechanical nonlinear coupling [6, 57]. Furthermore, generating largely delocalized mechanical states underpin the current effort to generate massive Schrodinger-cat states. Growing efforts have been made recently (see for example [5, 46]) to achieve this, as it provides a realistic pathw to both explore the macroscopic limit of quantum theory [35] and performing table-top experiments able to reveal coherent gravitational effects [7].
@@
In this work, we propose a protocol to herald coherent mechanical states upon post-selection of the emitted light on a photodetector in Geiger mode. Different experimental realizations of similar protocols have already been reported. In Ref. [13], an optical oscillator in wispering-galley-mode at GHz and room temperature is excited by continuous drive, and the entangled scattered component sent to a Single-Photon Avalanche Photodiod (SPAD). An analogue experiment using an optomechanical crystal is performed in [40], but with Superconducting Nanowire Spingle Photon Detector (SNSPD). Similarly, in Ref. [14], the authors realize a detection scheme that combines single-photon counting and optical heterodyne detection to engineer single-phonon addition and subtraction on wispering-gally mode. In Ref.[16] writing and reading driving pulses are shone on a 22ng membrane-in-the-middle oscillating at MHz frequency and cryogenic temperatures. Entangled modes are then detected by a SNSPD, generating nonclassical quantum mechanical state. Also in Ref. [17] a pulsed strategy is employed with a cavity optomechanical setup, showing single-photon state preparation of GHz mechanical oscillators at cryogenic temperatures. In order to resolve entangled optical modes, all these experiments operate in the sideband resolved regime.
@@
The protocol here proposed is effectively independent of the specific optomechanical platform, showing how cryogenic temperatures and an efficient collection of light emitted on the Stokes sideband are the sole crucial requirements to herald coherent mechanical states. We consider a $L=1$ mm long high finesse optomechanical cavity, were the reflective end oscillate at $\Omega=2\pi~10$ MHz with a mass of $m=50$ ng, and mechanical quality factor $Q=\Omega/\Gamma=10^{6}$. In section II, we discuss the optomechanical dynamics in: (i) a pulsed protocol driven by a squared blue-detuned beam, and (ii) a continuous-wave, steady-state protocol with further mode filtering. In section III the generation of optomechanical entanglement in the two case is discussed; we show that in the case of semicalssical input states (thermal), entanglement is mostly carried by photons emitted on the Stokes sideband. This phenomena is at the base of the results of section IV, where we show how in both cases it is possible to generate Wigner Functions (WFs) with negative regions, a criteria commonly considered as an indicator of nonclassicality [24, 33]. In general the pulsed case shows higher entanglement and heralding probabilities, while the steady state emission is as nonclassical as weak. Finally, in section V we test the robustness of this protocol to higher temperature. We include the temperature in the pulsed case by means of a Bogoliubov input-output treatment. The steady-state negativity shows great resilience to higher temperatures up to 2020K, while the mechanical coherences generated by pulses have high heralding probabilities.
@@ II.1 Linearization
In principle, radiation pressure coupling can generate coherent superpositions of mechanical states entangled with cavity Fock states [6], enabling the production of highly non-classical mechanical states (see for example Ref. [30]). However, in practice, both optical and mechanical dissipation rates typically exceed the bare coupling $g_{0}$, making these phenomena difficult to observe. It is therefore standard practice to enhance the effective optomechanical interaction by driving the system with strong input power, which boosts the coupling strength. This enhancement can be understood by examining the semiclassical equations of motion derived linearizing Eq. (1) for the expectation values $\langle a\rangle=\alpha$, $\langle q\rangle=Q$, and $\langle p\rangle=P$ under a mean-field approximation $\langle\hat{q}~\hat{a}\rangle\simeq Q\alpha$:
@@
Here, $\kappa$ and $\Gamma$ are the optical and mechanical dissipation rates accounting for photon leakage and Brownian noise from the thermal environment (for the linearization procedure see appendix A). Assuming these equations to be satisfied, the resulting linearized Hamiltonian describes the dynamics of the optical ($\hat{\delta a}$) and mechanical fluctuations ($\hat{\delta q},~\hat{\delta p}$) around their semiclassical values. It inherits its time dependence from the solutions of Eq. (2), through the detuning $\Delta(t)=\Delta_{0}-g_{0}Q(t)$ and the effective coupling $g(t)=g_{0}\alpha(t)$. Driving the cavity with strong input powers yields a highly populated intracavity field, thereby enhancing the interaction strength to $|\alpha|g_{0}$. This enhancement, however, comes at the cost of neglecting the intrinsic nonlinear optomechanical term $\sim O(g_{0})$. In what follows, we restrict to regimes where both $g(t)$ and $\Delta(t)$ can be regarded as approximately constant, and we denote them simply as $g$ and $\Delta$. Introducing the phononic modes opeartors $\hat{q}=(\hat{\delta b}+\hat{\delta b}^{\dagger})/\sqrt{2}$, $\hat{p}=i(\hat{\delta b}^{\dagger}-\hat{\delta b})/\sqrt{2}$, and moving into a rotating frame at frequency $\Delta$, the Hamiltonian
@@ II.3 Continuous drive
In the case of continuous driving, it is first necessary to examine the stability of the system’s dynamics. Optomechanical systems are known to display a rich variety of nonlinear behaviors, including multistable limit and unstable cycles [31, 59], routes to chaos [59, 4], and hysteresis or bistability [21]. In what follows we focus on the regime where the system admits a unique stable steady state. This usually verified for red-detuned drives ($\Delta>0$) and limited input power [43, 21]. In the stable regime, the balance between coherent drive and cavity losses establishes steady states characterized by the semiclassical averages
@@
The steady state output spectrum of the cavity varies depending on the input power. In the case where $\Omega\gg\kappa\gg g$, the sidebands are well resolved, and the RWA can still be applied. As steady state exist only for red-detuned drives, the output spectrum of the cavity will have the main peak at the carrier’s frequency $\omega_{l}$, and the resonant Anti-Stokes peak at $\omega_{l}+\Omega=\omega_{c}$. Stokes processes at $\omega_{l}-\Omega=\omega_{c}-2\Omega$ are instead off-resonant with the cavity, and for small $g/\Omega$ their bounded scattering amplitudes remain negligible (cf. appendix B). Things change when one increases the input power, i.e. the coupling $g$. As a first note, we remark that achieving the ultra-strong coupling regime $g\gg\kappa$ is challenging in a realistic setup [3] (although it has been observed in [22]), so in the following we restrict our analysis to the case where $g\simeq\kappa$.
@@ III Entanglement Generation
In our scheme, entanglement produced by the two-mode squeezing interaction is the fundamental resource that enables conditioning of the mechanical state on the outcome of an optical detection. Indeed, two-modes mixing proves to be very effective in state swapping protocols [23, 39, 55] or for sideband cooling [20, 11, 50], but cannot generate entanglement if the input state are semiclassical [26, 54].
@@ III.1 Pulsed regime
where only TMS resonant terms have been selected under RWA. In the limit $\kappa\gg g$, the cavity field can be adiabatically eliminated, setting $\dot{a}=0$, as its energy exchange are suppressed [51]. Formally, the dynamics are solved by
@@
$\hat{U}_{\text{pulse}}(r)=\hat{R}_{\hat{A}_{\text{in}}}(\pi)~\hat{S}\left(r,\frac{\pi}{2}\right)$ (13)
, where $\hat{R}_{\hat{A}_{\text{in}}}(\pi)=e^{-i\pi\hat{A}_{\text{in}}^{\dagger}\hat{A}_{\text{in}}}$ is a $\pi$-rotation in the phase space on the input optical field, and $\hat{S}(r,\frac{\pi}{2})=e^{-ir(\hat{A}_{\text{in}}^{\dagger}\hat{B}_{\text{in}}^{\dagger}-\hat{A}_{\text{in}}\hat{B}_{\text{in}})}$ a TMS transformation on the input modes with squeezing rate $r=\text{cosh}^{-1}(e^{G\tau})$. This is a well defined transformations, as it transforms the properly normalized input modes of the optical field $\hat{A}_{\text{in}},\hat{A}_{\text{in}}^{\dagger}]=1$ and mechanical modes $\hat{B}_{\text{in}},\hat{B}_{\text{in}}^{\dagger}]=1$, into output modes preserving canonical bosonic commutations relations $\hat{A}_{\text{out}},\hat{A}_{\text{out}}^{\dagger}]=1$, $\hat{B}_{\text{out}},\hat{B}_{\text{out}}^{\dagger}]=1$, and $\hat{A}_{\text{out}},\hat{B}_{\text{out}}^{\dagger}]=0 [$2, 18]. It is important to notice that now the optical modes $\hat{A}_{j}$ refer to the optical field of the environment, at the frequency $\Delta=-\Omega$. Therefore, one can define a stroboscopic evolution in the Schrödinger picture between the input and output optical-mechanical states
@@ III.2 Continuous drive
with $\hat{A}_{\text{out}},\hat{A}_{\text{out}}^{\dagger}]=\mathbb{1}$ only when $A_{\text{out}}=\underset{t\rightarrow\infty}{\text{lim}}A_{\text{out}}(t)$, as $\int_{-\infty}^{\infty}ds|\varphi(s)|^{2}=1$. Moving in the frequency domain by mean of the Fourier Transform $\mathcal{O}(t)=\int\frac{d\omega}{{2\pi}}\tilde{\mathcal{O}}(\omega)e^{-i\omega t}$, it results clear that the output mode
@@
is a $\text{sinc}(x)=\text{sin}(x)/x$ shaped filter. The definition of a bosonic mode allows the one of the output quadratures $\hat{X}_{\text{out}}=(\hat{A}_{\text{out}}+\hat{A}_{\text{out}}^{\dagger})/\sqrt{2}$, $\hat{Y}_{\text{out}}=i(\hat{A}_{\text{out}}^{\dagger}-\hat{A}_{\text{out}})/\sqrt{2}$. The input-output relation can then be written in the form
@@
Finally, establishing the CM  eq. 21 fully defines the state of the fluctuations around the semiclassical averages. In terms of gaussian states [15, 37], its WF can be readily expressed as
@@
When the mode has sufficient duration (bandwidth), it is able to capture entangled sibedands photons, effectively performing entanglement distillations, as pointed in fig. 1.
@@ IV Generation of delocalized mechanical states
The application of a projective measurememt on the emitted light introduce a nonlinear abruption of the dynamic, as described by the measurement postulate. Here, a photodetector in Geiger mode herald the detection, a click, or the absence, no-click, of a certain light mode. In this section we show that the event of a detection of entangled photons $\hat{A}_{\text{out}}$ from the Stokes sideband at $\omega_{l}-\Omega=\omega_{c}$ suffices to generate delocalized mechanical states. The POVMs for the event of no-click and click are
@@
$\hskip-56.9055pt\text{with}\hskip 56.9055ptp_{\text{click}}(\eta)=\big\langle\hat{\Pi}_{1}\big\rangle_{\rho_{\text{out}}}.$
@@
Figure 3: (a): WF of the conditional mechanical state for pulses duration of $\Omega\tau_{1}=10$ and $\Omega\tau_{2}=60$, fixed input power at $P=3$ mW and $n_{0}=0.1$. Both WF shows a negative region around the origin. The heralding probability is $\sim 10\%$ and $52\%$ respectively. In the first case, two-modes squeezing interaction excites only the first Fock states, whose WF shows to be more nonclassical than the one resulting from longer pulses, where higher Fock states are excited. Fig (b) illustrate this, showing relative probabilities of detection of each Fock state (blue for $\tau_{1}$ and orange for $\tau_{2}$).

### v2

@@ Abstract
Non-Gaussian mechanical states are a key resource for quantum-enhanced sensing and tests of macroscopic quantum physics. We propose a measurement-based protocol to herald delocalized, nonclassical states of a mechanical oscillator in cavity optomechanics by conditioning on Geiger photodetection of the optical output. We analyse under which conditions Stokes-induced optomechanical entanglement gives rise to mechanical Wigner Function negativity upon detection. We develop and compare a blue-detuned pulsed scheme and a continuous-wave steady-state scheme employing temporal-mode filtering, and we quantify heralding rates and robustness to finite temperature under realistic detection efficiencies.
Optomechanics, Non-Gaussian states, Non-Gaussian Measurements, Delocalized Mechanical States
Corresponding author: mbordin01@qub.ac.uk
@@ I Introduction
Quantum optomechanics offers a possible route to macroscopic quantum control: massive oscillators comprising billions of atoms can be cooled, entangled, and measured at or near the quantum limit. A challenge is the generation of delocalized coherent mechanical states. Previous strategies implemented in levitated and clamped-cavity platforms typically introduce a nonlinearity or conditional operation to evolve semiclassical inputs into nonclassical outputs; examples include transient potentials, photon addition or subtraction, and conditional measurements [47, 35, 31, 42, 26, 54, 14] or the intrinsic optomechanical nonlinear coupling [7, 59]. Furthermore, generating largely delocalized mechanical states underpin the current effort to generate massive Schrodinger-cat states. Growing efforts have been made recently (see for example [6, 48]) to achieve this, as it provides a realistic path to both explore the macroscopic limit of quantum theory [36] and performing table-top experiments able to reveal coherent gravitational effects [8].
@@
In this work, we propose a protocol to herald coherent mechanical states upon post-selection of the emitted light on a photodetector in Geiger mode. Different experimental realizations of similar protocols have already been reported. In Ref. [14], an optical oscillator in whispering-gallery-mode at GHz and room temperature is excited by continuous drive, and the entangled scattered component sent to a Single-Photon Avalanche Photodiod (SPAD). An analogue experiment using an optomechanical crystal is performed in [41], but with Superconducting Nanowire Single Photon Detector (SNSPD). Similarly, in Ref. [15], the authors realize a detection scheme that combines single-photon counting and optical heterodyne detection to engineer single-phonon addition and subtraction on whispering-gallery mode. In Ref.[17] writing and reading driving pulses are shone on a 22ng membrane-in-the-middle oscillating at MHz frequency and cryogenic temperatures. Entangled modes are then detected by a SNSPD, generating nonclassical quantum mechanical state. Also in Ref. [18] a pulsed strategy is employed with a cavity optomechanical setup, showing single-photon state preparation of GHz mechanical oscillators at cryogenic temperatures.  [^1] Note added in proof: The author acknowledges similar protocols have been recently proposed for levitated cavity optomechanics [5] and tested for superfluid Helium vibrations [45].
Figure 1: In (a): schematic representation of the cavity-optomechanical setup considered: a Fabry-Pérot cavity is driven by an external coherent drive. One end is free to oscillate, coupling phononic vibrations to photons. The output light is collected by a detector. (b) Typical output spectrum of an optomechanical cavity: the central lorentzian allows for sideband resolution in the sideband-resolved regime $\kappa\ll\Omega$. The sidebands correspond to quadratic inelastic scattering processes, shifting the energy of output photons of a mechanical quanta $\pm\hbar\Omega$. Depending on the cavity-drive detuning $\Delta_{0}=\omega_{c}-\omega_{l}$, sidebands can show strong asymmetry.
@@
The protocol here proposed is effectively independent of the specific optomechanical platform, showing how cryogenic temperatures and an efficient collection of light emitted on the Stokes sideband are the sole crucial requirements to herald coherent mechanical states. We consider a $L=1$ mm long high finesse optomechanical cavity, where the reflective end oscillates at $\Omega=2\pi~10$ MHz with a mass $m=50$ ng and mechanical quality factor $Q=\Omega/\Gamma=10^{6}$. In section II, we discuss the optomechanical dynamics in: (i) a pulsed protocol driven by a squared blue-detuned beam, and (ii) a continuous-wave, steady-state protocol with further mode filtering. In section III the generation of optomechanical entanglement in the two case is discussed; we show that in the case of semiclassical input states (thermal), entanglement is mostly carried by photons emitted on the Stokes sideband. This phenomena is at the base of the results of section IV, where we show how in both cases it is possible to generate Wigner Functions (WFs) with negative regions, a criteria commonly considered as an indicator of nonclassicality [25, 34]. In general the pulsed case shows higher entanglement and detection probabilities, while the steady state emission herald strongly nonclassical states due to its strong nonclassicality. Finally, in section V we test the robustness of this protocol to higher temperature. We include the temperature in the pulsed case by means of a Bogoliubov input-output treatment. The steady-state negativity shows great resilience to higher temperatures up to 2020K, while the mechanical coherences generated by pulses have high heralding probabilities.
@@ II.1 Linearization
In principle, radiation pressure coupling can generate coherent superpositions of mechanical states entangled with cavity Fock states [7], enabling the production of highly non-classical mechanical states (see for example Ref. [31]). However, in practice, both optical and mechanical dissipation rates typically exceed the bare coupling $g_{0}$, making these phenomena difficult to observe. It is therefore standard practice to enhance the effective optomechanical interaction by driving the system with strong input power, which boosts the coupling strength. This enhancement can be understood by examining the semiclassical equations of motion, derived by linearizing Eq. (1) for the expectation values $\langle a\rangle=\alpha$, $\langle q\rangle=Q$, and $\langle p\rangle=P$ under a mean-field approximation $\langle\hat{q}~\hat{a}\rangle\simeq Q\alpha$:
@@
Here, $\kappa$ and $\Gamma$ are the optical and mechanical dissipation rates accounting for photon leakage and Brownian noise from the thermal environment (for the linearization procedure see appendix A). Assuming these equations to be satisfied, the resulting linearized Hamiltonian describes the dynamics of the optical ($\hat{\delta a}$) and mechanical fluctuations ($\hat{\delta q},~\hat{\delta p}$) around their semiclassical values. It inherits its time dependence from the solutions of Eq. (2), through the detuning $\Delta(t)=\Delta_{0}-g_{0}Q(t)$ and the effective coupling $g(t)=g_{0}\alpha(t)$. Driving the cavity with strong input powers yields a highly populated intracavity field, thereby enhancing the interaction strength to $|\alpha|g_{0}$. This enhancement, however, comes at the cost of neglecting the intrinsic nonlinear optomechanical term $\sim O(g_{0})$. In what follows, we restrict to regimes where both $g(t)$ and $\Delta(t)$ can be regarded as approximately constant, and we denote them simply as $g$ and $\Delta$. Introducing the phononic mode operators $\hat{q}=(\hat{\delta b}+\hat{\delta b}^{\dagger})/\sqrt{2}$, $\hat{p}=i(\hat{\delta b}^{\dagger}-\hat{\delta b})/\sqrt{2}$, and moving into a rotating frame at frequency $\Delta$, the Hamiltonian
@@ II.3 Continuous drive
In the case of continuous driving, it is first necessary to examine the stability of the system’s dynamics. Optomechanical systems are known to display a rich variety of nonlinear behaviors, including multistable limit and unstable cycles [32, 61], routes to chaos [61, 4], and hysteresis or bistability [22]. In what follows we focus on the regime where the system admits a unique stable steady state. This is usually verified for red-detuned drives ($\Delta>0$) and limited input power [44, 22]. In the stable regime, the balance between coherent drive and cavity losses establishes steady states characterized by the semiclassical averages
@@
The steady state output spectrum of the cavity varies depending on the input power. In the case where $\Omega\gg\kappa\gg g$, the sidebands are well resolved, and the RWA can still be applied. As steady state exist only for red-detuned drives, the output spectrum of the cavity will have the main peak at the carrier’s frequency $\omega_{l}$, and the resonant Anti-Stokes peak at $\omega_{l}+\Omega=\omega_{c}$. Stokes processes at $\omega_{l}-\Omega=\omega_{c}-2\Omega$ are instead off-resonant with the cavity, and for small $g/\Omega$ their bounded scattering amplitudes remain negligible (cf. appendix B). Different effects arises when one increases the input power, i.e. the coupling $g$. As a first note, we remark that achieving the ultra-strong coupling regime $g\gg\kappa$ is challenging in a realistic setup [3] (although it has been observed in [23]), so in the following we restrict our analysis to the case where $g\simeq\kappa$.
@@ III Entanglement Generation
In our scheme, entanglement produced by the two-mode squeezing interaction is the fundamental resource that enables conditioning of the mechanical state on the outcome of an optical detection. Indeed, two-modes mixing proves to be very effective in state swapping protocols [24, 40, 57] or for sideband cooling [21, 12, 52], but cannot generate entanglement if the input state is semiclassical [27, 56].
@@ III.1 Pulsed regime
where only TMS resonant terms have been selected under RWA. In the limit $\kappa\gg g$, the cavity field can be adiabatically eliminated, setting $\dot{a}=0$, as its energy exchange is suppressed [53]. Formally, the dynamics are solved by
@@
$\hat{U}_{\text{pulse}}(r)=\hat{R}_{\hat{A}_{\text{in}}}(\pi)~\hat{S}\left(r,\frac{\pi}{2}\right),$ (13)
where $\hat{R}_{\hat{A}_{\text{in}}}(\pi)=e^{-i\pi\hat{A}_{\text{in}}^{\dagger}\hat{A}_{\text{in}}}$ is a $\pi$-rotation in the phase space on the input optical field, and $\hat{S}(r,\frac{\pi}{2})=e^{-ir(\hat{A}_{\text{in}}^{\dagger}\hat{B}_{\text{in}}^{\dagger}-\hat{A}_{\text{in}}\hat{B}_{\text{in}})}$ a TMS transformation on the input modes with squeezing rate $r=\text{cosh}^{-1}(e^{G\tau})$. This is a well defined transformation, as it transforms the properly normalized input modes of the optical field $\hat{A}_{\text{in}},\hat{A}_{\text{in}}^{\dagger}]=1$ and mechanical modes $\hat{B}_{\text{in}},\hat{B}_{\text{in}}^{\dagger}]=1$, into output modes preserving canonical bosonic commutations relations $\hat{A}_{\text{out}},\hat{A}_{\text{out}}^{\dagger}]=1$, $\hat{B}_{\text{out}},\hat{B}_{\text{out}}^{\dagger}]=1$, and $\hat{A}_{\text{out}},\hat{B}_{\text{out}}^{\dagger}]=0 [$2, 19]. It is important to notice that now the optical modes $\hat{A}_{j}$ refer to the optical field of the environment, at the frequency $\Delta=-\Omega$. Therefore, one can define a stroboscopic evolution in the Schrödinger picture between the input and output optical-mechanical states
@@ III.2 Continuous drive
with $\hat{A}_{\text{out}},\hat{A}_{\text{out}}^{\dagger}]=\mathbb{1}$ only when $A_{\text{out}}=\underset{t\rightarrow\infty}{\text{lim}}A_{\text{out}}(t)$, as $\int_{-\infty}^{\infty}ds|\varphi(s)|^{2}=1$. Moving in the frequency domain by mean of the Fourier Transform $\mathcal{O}(t)=\int\frac{d\omega}{{2\pi}}\tilde{\mathcal{O}}(\omega)e^{-i\omega t}$, it results that the output mode
@@
is a $\text{sinc}(x)=\text{sin}(x)/x$ shaped filter. The definition of a bosonic mode allows output quadratures to be defined as $\hat{X}_{\text{out}}=(\hat{A}_{\text{out}}+\hat{A}_{\text{out}}^{\dagger})/\sqrt{2}$, $\hat{Y}_{\text{out}}=i(\hat{A}_{\text{out}}^{\dagger}-\hat{A}_{\text{out}})/\sqrt{2}$. The input-output relation can then be written in the form
@@
Finally, establishing the CM  eq. 21 fully defines the state of the fluctuations around the semiclassical averages. A useful set of results from the theory of Gaussian states [16, 38] can be then be applied: it is well established that the WF associated to a Gaussian state can be readily expressed as
@@
When the mode has sufficient duration (bandwidth), it is able to capture entangled sideband photons, effectively performing entanglement distillations, as shown in fig. 2.
@@ IV Generation of delocalized mechanical states
The application of a projective measurement on the emitted light introduces a nonlinear abruption of the dynamic, as described by the measurement postulate. Here, a photodetector in Geiger mode heralds the detection, a click, or the absence, no-click, of a certain light mode. In this section we show that the event of a detection of entangled photons $\hat{A}_{\text{out}}$ from the Stokes sideband at $\omega_{l}-\Omega=\omega_{c}$ suffices to generate delocalized mechanical states. The POVMs for the event of no-click and click are
@@
$\hskip-56.9055pt\text{with}\hskip 56.9055ptp_{\text{click}}(\eta)=\big\langle\hat{\Pi}_{1}\big\rangle_{\rho_{\text{out}}}$
@@
Figure 4: (a): WF of the conditional mechanical state for pulses duration of $\Omega\tau_{1}=10$ and $\Omega\tau_{2}=60$, fixed input power at $P=3$ mW and $n_{0}=0.1$. Both WF shows a negative region around the origin. The heralding probability is $\sim 10\%$ and $52\%$ respectively. In the first case, two-modes squeezing interaction excites only the first Fock state, whose WF shows to be more nonclassical than the one resulting from longer pulses, where higher Fock states are excited. Fig (b) illustrate this, showing relative probabilities of detection of each Fock state (blue for $\tau_{1}$ and orange for $\tau_{2}$).